Set up a sum of integrals that will give the area of the quadrilateral with vertices at (0,0), (1,1), (0,2), and (2,1), connected in the stated order.



1
Expert's answer
2021-02-02T04:38:52-0500



"1. BC\implies y-2=\frac{2-1}{0-2}(x-0)\implies y-2=\frac{1}{2}(x)"

"BC\implies 2y+4=x\implies x=4-2y"


"2. CA\implies y-1=\frac{1-0}{2-0}(x-2)\implies y-1=\frac{1}{2}(x-2)"

"CA\implies 2y-2=x-2\implies x=2y"


"3. BD\implies y-2=\frac{2-1}{0-1}(x-0)\implies y-2=\frac{1}{-1}(x)"

"BD\implies -y+2=x\implies x=2-y"


"4. DA\implies y-1=\frac{1-0}{1-0}(x-1)\implies y-1=\frac{1}{1}(x-1)"

"DA\implies y-1=x-1\implies x=y"


Area of shaded region ="\int_1^2x_{BC}dy-\int_1^2x_{BD}dy+\int_0^1x_{CA}dy-\int_0^1x_{DA}dy"

"\int_1^2(4-2y)dy-\int_1^2(2-y)dy+\int_0^12ydy-\int_0^1ydy"

"[4y-y^2]_1^2-[2y-0.5y^2]_1^2+[y^2]_0^1-[0.5y^2]_0^1"

"1-0.5+0.5+0"

"1"